OpenStudy (anonymous):

hi i need help with this problem 20=-3/8d

OpenStudy (anonymous):

d=-3/160

OpenStudy (anonymous):

\[20 = - \frac{3}{8d}\]\[(20)(8d) = -3\]

OpenStudy (anonymous):

\[160d = -3\]\[d = - \frac{3}{160}\]